If a boy lifts a mass of 6kg to a height of 10m and travels horizontally with a constant velocity of 4.2m/s, calculate the work done? Explain your answer.


1
Expert's answer
2021-07-21T09:05:46-0400

The work done is zero. In this case the force acts upward on the mass as it is displaced rightward. As a result, the force vector and the displacement vector are at right angles to each other. Therefore, the angle between the force and displacement is "90^{\circ}" and we get:


"W=Fdcos\theta=mghcos\theta=6\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot10\ m\cdot cos90^{\circ}=0."